OpenStudy (anonymous):

I'm finding this one difficult. 3/2(x-1) - x+2|3 = 1 - x +11|6 the part with x + 11 |6 means x+11 divided by 6 same with x+2|3 means x+2 divided by 3, I didnt know how to write that. Can someone please show me the steps to this it would be greatly appreciated ! :)

OpenStudy (anonymous):

You asked for it.

OpenStudy (anonymous):

(3)/(2)*(x-1)-x+(2)/(3)=1-x+(11)/(6) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 3. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (3(x-1))/(2)-x*(3)/(3)+(2)/(3)=1-x+(11)/(6) Complete the multiplication to produce a denominator of 3 in each expression. (3(x-1))/(2)-(3x)/(3)+(2)/(3)=1-x+(11)/(6) Combine the numerators of all expressions that have common denominators. (3(x-1))/(2)+(-3x+2)/(3)=1-x+(11)/(6) Multiply each term by a factor of 1 that will equate all the denominators. In this case, all terms need a denominator of 6. The (3(x-1))/(2) expression needs to be multiplied by ((3))/((3)) to make the denominator 6. The ((-3x+2))/(3) expression needs to be multiplied by ((2))/((2)) to make the denominator 6. (3(x-1))/(2)*(3)/(3)+(-3x+2)/(3)*(2)/(2)=1-x+(11)/(6) Multiply the expression by a factor of 1 to create the least common denominator (LCD) of 6. (3(x-1)(3))/(6)+(-3x+2)/(3)*(2)/(2)=1-x+(11)/(6) Multiply 3*3 to get 9 in the numerator. ((9)(x-1))/(6)+(-3x+2)/(3)*(2)/(2)=1-x+(11)/(6) Remove the single term factors from the expression. (9(x-1))/(6)+(-3x+2)/(3)*(2)/(2)=1-x+(11)/(6) Multiply 9 by each term inside the parentheses. (9x-9)/(6)+(-3x+2)/(3)*(2)/(2)=1-x+(11)/(6) Multiply the expression by a factor of 1 to create the least common denominator (LCD) of 6. (9x-9)/(6)+((-3x+2)(2))/(6)=1-x+(11)/(6) Remove the parentheses in the numerator and move the single term to the front of the expression. (9x-9)/(6)+(2(-3x+2))/(6)=1-x+(11)/(6) Multiply 2 by each term inside the parentheses. (9x-9)/(6)+(-6x+4)/(6)=1-x+(11)/(6) The numerators of expressions that have equal denominators can be combined. In this case, ((9x-9))/(6) and ((-6x+4))/(6) have the same denominator of 6, so the numerators can be combined. ((9x-9)+(-6x+4))/(6)=1-x+(11)/(6) Simplify the numerator of the expression. (9x-9-6x+4)/(6)=1-x+(11)/(6) Combine all similar terms in the polynomial 9x-9-6x+4. (3x-5)/(6)=1-x+(11)/(6) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 6. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (3x-5)/(6)=-x+1*(6)/(6)+(11)/(6) Complete the multiplication to produce a denominator of 6 in each expression. (3x-5)/(6)=-x+(6)/(6)+(11)/(6) Combine the numerators of all fractions that have common denominators. (3x-5)/(6)=-x+(6+11)/(6) Add 11 to 6 to get 17. (3x-5)/(6)=-x+(17)/(6) Multiply each term in the equation by 6. (3x-5)/(6)*6=-x*6+(17)/(6)*6 Simplify the left-hand side of the equation by canceling the common factors. 3x-5=-x*6+(17)/(6)*6 Simplify the right-hand side of the equation by simplifying each term. 3x-5=-6x+17 Since -6x contains the variable to solve for, move it to the left-hand side of the equation by adding 6x to both sides. 3x-5+6x=17 Since 3x and 6x are like terms, add 6x to 3x to get 9x. 9x-5=17 Since -5 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 5 to both sides. 9x=5+17 Add 17 to 5 to get 22. 9x=22 Divide each term in the equation by 9. (9x)/(9)=(22)/(9) Simplify the left-hand side of the equation by canceling the common factors. x=(22)/(9)

OpenStudy (anonymous):

woah omg I love you thanks so much !!!!!!!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

If I'm reading your original equation correctly, I got x = 1.

OpenStudy (anonymous):

actually ya dmancine your right I just checked the answer, could you possibly tell me how you got it ?

OpenStudy (anonymous):

First, I'm going to use the Equation editor to make your equation easier to read.\[\frac{3}{2}(x-1)-\frac{x+2}{3}=1-\frac{x-11}{6}\]Is that the actual problem, or did I misunderstand?

OpenStudy (anonymous):

Ya you got it right, thats the actual problem

OpenStudy (anonymous):

The first thing I would do is get rid of the denominators by multiplying both sides by the same number. What number should we use?

OpenStudy (anonymous):

Alright , I used 6 which was the lowest common multiple.

OpenStudy (anonymous):

also the last part is = 1 - x + 11 over 6

OpenStudy (anonymous):

You should use parentheses to group things together so it's not ambiguous. Do you mean (1-x+11)/6 or 1-((x+11)/6)? Or did I just flip a sign when I wrote the equation before? (It looks like I just flipped the sign. Using x+11 yields x=1, which you said is right.) Order of operations is important. Multiplication and division happen before addition and subtraction. If you type x+11/6 that means \[x+\frac{11}{6}\]If you mean\[\frac{x+11}{6}\]you need to use parentheses to override the normal order of operations: (x+11)/6. Mathematics has a very precise language for expressing these things, and if we all use it we can communicate unambiguously. Just so we're clear, is this the problem: \[\frac{3}{2}(x-1) -\frac{x+2}{3}=1-\frac{x+11}{6}\] If so, post what you got just after you multiplied by 6.

OpenStudy (anonymous):

sorry it was (x+11)/6 and I have 9x-9-2x+2=1-x+6

OpenStudy (anonymous):

I think you've done more than one step there. Just multiply by 6 and don't distribute into the parentheses.

OpenStudy (anonymous):

ya i think i did so it would be 9(x-1)-2x+2=1-x +6

OpenStudy (anonymous):

Let's go one term at a time. What do you get when you multiply each of these terms by 6:\[\frac{3}{2}(x-1)\]\[-\frac{x+2}{3}\]\[1\]\[-\frac{x+11}{6}\]Just do the multiplication. Don't distribute. (You got the first one right, by the way.)

OpenStudy (anonymous):

alright awesome, i got 9(x-1) - (6x+2)/3 6 - (6x + 66)/6

OpenStudy (anonymous):

I think that second one is a little wrong still. Also, you're still distributing the 6 into the parentheses. Can you leave it out and leave the (x+2) and (x+11) parts the way they are?

OpenStudy (anonymous):

(I guess I might be confusing you by saying "into the parentheses" when there aren't actually any parentheses there. There's kinda "implied" parentheses in the 2nd and 4th terms because when you write them like -(x+2)/3 the parentheses show up.)

OpenStudy (anonymous):

ya i kinda got that part although its still a little confusing but would I just be multiplying the denominator by 6 thenand leaving the top part out ?

OpenStudy (anonymous):

Yes. The term\[-\frac{x+2}{3}\]really means the same thing as \[-\frac{1}{3}(x+2)\]if that helps. Do you understand why?

OpenStudy (anonymous):

uhm that helps but i dont understand

OpenStudy (anonymous):

oh does the (-) sign equal 1 and you just multiply (x+2)

OpenStudy (anonymous):

Well, taking the negative of something is the same as multiplying by -1. Also, dividing by 3 is the same as multiplying by 1/3. Do you understand both of those concepts?

OpenStudy (anonymous):

Okay ya I get that.

OpenStudy (anonymous):

Also, when x+2 is in the numerator you're dividing the whole quantity x+2 by the denominator.

OpenStudy (anonymous):

Uhm can you show me like a step I dont get that Im more of a visual learner.

OpenStudy (anonymous):

Does this help?\[-\frac{x+2}{3} = -\frac{(x+2)}{3}\]

OpenStudy (anonymous):

uhm whats it become ?

OpenStudy (anonymous):

oh nevermind I got that.

OpenStudy (anonymous):

So, before we do anything else, can you write the term\[-\frac{x+11}{6}\]as a fraction times (x+11), like I did above for the other term?

OpenStudy (anonymous):

like this ? -((x+11))/6

OpenStudy (anonymous):

oops (x+11)/6

OpenStudy (anonymous):

I was talking about when I did this:\[-\frac{x+2}{3} = -\frac{1}{3}(x+2)\]See the fraction out front?

OpenStudy (anonymous):

oh would it be ? 1/6 (x+11)

OpenStudy (anonymous):

Yes, except you lost the minus sign.

OpenStudy (anonymous):

oh oops forgot about that

OpenStudy (anonymous):

ok so its - 1/6 (x+11)

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

ok after these steps I keep distributing the 6 and I keep getting different answers

OpenStudy (anonymous):

Ok. Go back to the post where I had the 4 terms listed by themselves, and show me what you get for each of them when you just multiply them by 6. Don't distribute.

OpenStudy (anonymous):

ok I have 6(x-1) -2(x+2) 6 -6(x+11)

OpenStudy (anonymous):

The first one's wrong, but you did it right the first time (many posts ago). And the last one's wrong. Tell me how you got what you got for them.

OpenStudy (anonymous):

oh the first ones 9(x-1) i just messed that up and the seconds -1(x+11)

OpenStudy (anonymous):

Yes! Okay. Now we have 9(x-1) -2(x+2) 6 -1(x+11) For all but the third one, distribute the number into the parentheses.

OpenStudy (anonymous):

awesome so 9x-1 -2x+(-4) 6 -1x +(-11)

OpenStudy (anonymous):

First one's a little wrong.

OpenStudy (anonymous):

oh 9x - 9 sorry Im really tired

OpenStudy (anonymous):

lol. I understand. And you did the other ones right, so I figure you know what you're doing with distributing. But we have to get every little step right. Okay. It'll be easier to deal with if you change the things like +(-4) into just a subtraction. And -1x is just -x. So, do those steps.

OpenStudy (anonymous):

ya thats true its helping alot. so it would be 9x - 1 -2x-4 6 -x-11

OpenStudy (anonymous):

I'll assume the first was just a typo again. Now, we need to assemble those pieces back into the equation we took them out of. Try doing that.

OpenStudy (anonymous):

I'll assume the first was just a typo again. Now, we need to assemble those pieces back into the equation we took them out of. Try doing that.

OpenStudy (anonymous):

i think I got it 9x - 9 - 2x-4=6 - x-11 9x+x-2x=6-11+9+4 8x=8 x=8/8 x=1

OpenStudy (anonymous):

Yep. But make sure you understand how we went from the original problem to the equation we got after multiplying by 6 and distributing.